Calculate the amount of oxygen gas (O) in a cylinder of 30L,if the pressure is 20atm at 30 kelvin







1
Expert's answer
2022-01-25T14:46:02-0500

according to ideal gas low, PV = nRT

Where,

P= pressure

V = volume

n = mole

R = avogedro`s constant = 8.314 J/mol K

T = Temperature

in this problem V=30L = 0.030 m3,

P=20 atm = 20 × 101325 Kpa,

T=30C = 303 K

So, n = PV/RT

= (20×101325) × (0.030) / (8.314) × (303)

= 24.13 mol (for oxygen 1 mol O2 = 32 gm)

= 772. 26 gm

So, the amount of oxygen in the cylinder is 772.26 gm